# Multi-scale phylodynamic modelling of rapid punctuated pathogen evolution

### v2

@@ 4.2 Multi-scale phylodynamic simulation
This section describes processing layers (A), (C) and (D), highlighting their key dynamic relationships. Since the demographic characteristics of agents, including their social groups, are not updated during simulation, there is no dynamic modelling involved in layer (B), and this layer is described in Supplementary Section B.
@@ 4.2.1 Phylogenetic layer (A)
The phylogenetic layer models the within-host evolution (mutation and selection), using pathogen genomes associated with each infected agent, and computes the fitness of circulating variants. An artificial genome contains 3,090 nucleotides (or 1,030 codons) which are associated with 20 known types of amino acids. The genome structure is partitioned into the spike region (100 codons) and the non-spike region (930 codons), with an overlapping region accounting for epitopes (75 codons, with 45 of these located in the spike region and 30 in the non-spike region). Supplementary Section A.1 provides a detailed description of the genome structure, followed by Supplementary Section A.2 describing point mutations.
The pathogen fitness is defined in terms of its transmissibility, reflected in the corresponding basic reproductive number ($R_{0}$). To quantify the individual fitness contributions of amino acids at each codon position, we employ a weight table, specifying weights $a_{i,j}$ of $N_{A}=20$ types of amino acids across $N=1,030$ codon positions, as detailed in Supplementary Section A.3. The overall fitness $K$ of a strain with genome $s$ is determined as the sum of these individual contributions:
@@
To re-iterate, the phylogenetic model distinguishes between two infected host categories: typical infected hosts and chronically infected hosts (Supplementary Section A.4). Consequently, there are different selective pressures driving the within-host evolution of pathogens in these categories (Supplementary Section A.5).
We distinguish between mutation and substitution rates. While the mutation rate refers to the frequency of new mutations arising in a genome per unit of time, the substitution rate is the rate at which these new mutations are retained over time within the population [33, 34, 31]. In our case study, the mutation rate is a key input parameter, whereas the average substitution rate emerges as a simulation outcome. We use the attained substitution rate to calibrate the model, by matching the regression coefficient observed in empirical observations, illustrated in Fig. 2 (B).
@@ 4.2.2 Epidemic layer (C)
As described in Supplementary Section C.1, disease is transmitted among agents that interact across different social contexts. The transmission process is simulated in discrete half-day time steps: “daytime” cycles during which agents interact in workplaces or educational settings (e.g., class, grade, school), and “nighttime” cycles during which agents interact in residential settings (e.g., household, household cluster, neighbourhood, and community). Each agent progresses through several health states: Susceptible, Infectious (asymptomatic or symptomatic), and Recovered, following the natural history of the disease.
At simulation cycle $n$, the infection probability $p_{i}(n)$ for a susceptible agent $i$ is determined across all its social contexts $g\in G_{i}$ (see Supplementary Section C.1.1):
@@
where $q_{j\rightarrow i}(g)$ is the age-dependent interaction probability within $g$ (see Supplementary Section C.1.1); $n_{j}$ denotes the infection onset time for agent $j$; the agent-specific function $f_{j}(n-n_{j})$ is the natural history of the disease, reflecting the infectivity of agent $j$ as its infection progresses (see Supplementary Section C.1.2); and $K(s_{j})$ represents the transmissibility of pathogen variant $s_{j}$, proportional to the corresponding basic reproductive number $R_{0}$, i.e., $K(s_{j})$ is the fitness of genome $s_{j}$ carried by agent $j$, as defined by equation (1).
Once susceptible agent $i$ becomes infected, it is possible to assign (i.e., identify) a specific infectious agent as the source of infection. This is simulated by weighted randomly sampling an infection source from all potential infectious agents $j$ across all social contexts $g\in G_{i}(n)$ in which agent $i$ interacted during this cycle. Then the pathogen genome profile $s_{j}$ (carried by the identified infection source agent $j$) is copied to agent $i$ (see Supplementary Section C.1.1).
Various interventions may change the infection probabilities across social contexts. Modelling non-pharmaceutical interventions (NPIs) is described in Section C.2.
@@ 4.2.3 Immunological layer (D)
The immunological layer simulates a vaccination rollout, including vaccination coverage, schedule and rates, given vaccine efficacy. Furthermore, the immuno-epidemiological model quantifies the agent immunity resulting from multiple immunity-boosting events (i.e., vaccinations and infections), as described in Supplementary Section D.1. In this work, we broaden the concept of “hybrid immunity” to “compound immunity”, in order to capture non-linear immunity accumulation over various combinations of prior vaccinations and infections. Compound immunity may result from one or multiple vaccinations, one or multiple past infections, or a combination of both vaccination(s) and past infection(s) [27].
@@
where $m_{i}^{\textrm{c}}(n,r,s_{x})$ is the immunity against symptomatic infection induced by past immunological event $r$ (vaccination or infection); $H_{i}$ is the immunological history formed by past records $r$ up to cycle $n$; and $s_{x}$ is the genome carried by infectious agent $x$. In addition, the compound immunity wanes over time and depends on the genetic distance (see equation (12) in Supplementary Section D.2.1). Supplementary Section D.2 provides more details accounting for different vaccine efficacy components contributing to the compound immunity, and its effects on infection probabilities.
@@ 4.3.1 Hamming distance
Figures 2 and 3 trace the reference and pairwise Hamming distances for actual SARS-CoV-2 sequence data [15].
In order to trace Hamming distances for the simulated dynamics we followed slightly altered workflows, without applying the weekly windows for genome selection. When dealing with actual sequence data, these windows were needed to filter out sampling inconsistencies. Simulated data include pathogen genomes from all detected hosts, sampled on each simulation day, thus reducing sampling inconsistencies. The reference Hamming distance $\widehat{D}$ was computed against the simulated ancestral strain, constructed for each realisation (as described in Section 4). Figures 7, 911 trace the reference and pairwise Hamming distances for the simulated phylodynamics.
@@ 4.3.2 Statistical stationarity tests
ADF test. An ADF test detects the presence of unit root [37]. We use the following null and alternative hypotheses to determine stationary:
@@
We computed p-value from the ADF test and compared it against a chosen significance level (i.e., 0.05), with p-value smaller than the significance level rejecting $H_{0}$ and confirming stationarity. Results of ADF are shown in Supplementary section F.2 and Fig. 27.
CUSUM. A one-sided Cumulative Sum (CUSUM) control chart  [38, 39] can identify anomalies in the observed time series. In this study, we applied CUSUM on pairwise Hamming distance as shown in Fig. 10.
@@
The high and low CUSUM values, $S^{X}$ and $S^{Y}$, are traced in Fig. 10. We then applied a positive peak detection in $S^{X}$ and $S^{Y}$ (i.e., considering CUSUM value greater than 0), to detect the anomalies. Specifically, the peaks identified in $S^{X}$ indicate the emergence of a more transmissible variant, whereas the peaks identified in $S^{Y}$ indicate the dominance of a variant within the population. The number of dominant variants is equivalent to the total number of peaks detected in both $S^{X}$ and $S^{Y}$.
@@ Acknowledgement
Supplementary Materials also include some additional results (Section E), sensitivity analysis (Section G) and details on the computational complexity and implementation of PhASE TraCE (Section H).
@@ Appendix A Phylogenetic model
  • Genome structure represented by a nucleotide sequence with positions assigned to different functions (Section A.1).
  • Point mutations occurring at a rate that is averaged across the genome (Section A.2).
  • Pathogen fitness determined based on amino acids at specific regions of the genome(Section A.3).
  • Infected host categories, differentiating between typical infected hosts and chronically infected hosts (Section A.4).
  • Within-host (or intra-host) selective pressure generating advantageous pathogen mutations (Section A.5).
@@ A.1 Genome structure
Each nucleotide position in the simulated genome is represented by a number between 0 and 3. Every three nucleotides is interpreted as a codon (3,090 nucleotides correspond to 1,030 codons, see Fig. 13), encoding a sequence of amino acids following the standard translation rules [40]. We annotate the 20 known amino acids by numerical labels from 0 to 19 instead of the conventional letters [40]. Although we model translation to an amino acid sequence, we do not look for coding regions or otherwise interpret or annotate the simulated sequence in the context of the real virus.
@@
The simulated genome is partitioned into two distinct regions (Fig. 13): the “spike” region (consisting of 100 codons) and the non-spike region (consisting of 930 codons). This partitioning affects only two aspects of the simulation: (1) vaccine-related selective pressure is restricted to the spike region, and (2) mutations in the spike region have a greater chance of enhancing transmissibility (Supplementary Section A.3). 75 codons (45 codons in the spike region and 30 codons in the non-spike region) are interpreted as epitopes relevant for modelling of immunity [41, 42]. 1,000 codons (including 45 that are also in epitopes) are interpreted as relevant to transmissibility. Section D.2.3 provides more details.
@@ A.2 Mutations
In each half-day simulation cycle, the total number of mutations ($L$) in each of the viral genomes within infected hosts is sampled from a Poisson distribution. The mean of the distribution is set based on an average mutation rate of 0.001 per nt per year across the genome, i.e., $3090\times 0.001/(365\times 2)\approx 0.00423$. The average mutation rate was chosen within the reported range for SARS-CoV-2 of 0.000219 to 0.0012 per nt per year [43, 31]. The relationship between mutations and their contribution to fitness is described in Supplementary Section A.3.
@@ A.3 Contribution to fitness
Individual mutations in the genome affect the overall viral fitness, resulting in growing transmissibility and higher reproductive number ($R_{0}$) [16, 44, 45]. Obermeyer et al. [16] assessed the relative fitness of SARS-CoV-2 lineages (measured as the fold increase in $R$ over the ancestral strain) by linearly combining fitness contributions of individual amino acid substitutions, some of which were found to be fitness-increasing and recognised as spike mutations and non-spike mutations within the nucleocapsid and nonstructural proteins. Additionally, Thadani et al. [45] identified a general distribution of amino acid selection at each position in SARS-CoV-2 sequences and highlighted positions that are prone to viral escape and mutations.
@@
Figure 14 shows a section of the weight table (the first 26 codon positions). The weight table generated for a simulation regulates the pathogen’s evolution. In the absence of other pressures, one would expect substitutions that increase fitness to be preferentially fixed as a simulation progresses, leading to a gradual increase in transmissibility. Using the weight table, we calculate the overall fitness $K$ of genome $s$ by combining the individual contributions of the present amino acids:
@@
where $a_{i,s(i)}$ is the fitness contribution of amino acid $j=s(i)$ located at codon position $i$, with $j\in[0,N_{A}-1]$. We note that $a_{i,s(i)}$ can be zero, positive, or negative, representing mutations of neutral, advantageous, or disadvantageous contribution to pathogen fitness.
@@
Prior to simulating the COVID-19 case study, we generated a genome for the ancestral SARS-CoV-2 strain sampled such that its viral fitness, calculated using Eq. 7 and the weight table, fell between 2.65 and 2.85, matching the transmissibility range of the ancestral SARS-CoV-2 strain modelled in our previous work [6]. We used this ancestral genome as a starting point from which other variants originate, thus assuming a shared ancestry for all variants [48, 49]. To optimise memory usage, given the high number of simulated infections, each strain is represented by a list of mutations against the ancestral genome instead of a full-length genome.
@@ A.4 Infected host categories
Host factors, including individual variation in immune response, are known to strongly affect epidemic progression [50, 51]. In particular, chronic carriage of SARS-CoV-2 by immunocompromised hosts is thought to have had a key role in the emergence of new viral variants during the pandemic [31]. In a host with compromised or suppressed immune function, the virus is subjected to different selective pressures and within-host dynamics. A host immune system that is not able to effectively clear the virus is likely to result in a longer infection duration, a higher viral load within the host, and a greater diversity of viral sub-populations [31]. These conditions may facilitate more substitutions and larger increases to viral fitness (saltational evolution) compared to evolution of the virus over the same interval of time when transmitted by a chain of typical hosts and subjected to transmission bottlenecks and competent immune systems. In the case of SARS-CoV-2, chronically infected hosts can have a prolonged recovery period of several months, compared to one to two weeks for typical infected hosts [52, 46].
To examine the possible role of chronic infections on the pathogen evolution, we differentiate between two infected host categories: chronically infected hosts and the typical infected hosts. The main distinctions include the following aspects: (i) chronically infected hosts are assigned a different natural history model with a significantly longer recovery period, randomly sampled from a uniform distribution ranging from 60 to 370 days reported in many studies (Fig. 15)[46, 53, 54], and (ii) chronically infected hosts may have a higher intra-host selective pressure (detailed in Section 29).
Ghafari et al. [46] reported that chronically infected hosts make up a small fraction of the infected hosts (between 0.1 and 0.5% of COVID-19 infections), and only 32% of these hosts may experience strong positive intra-host selection. Considering the high attack rate of the COVID-19, we assume that 0.1% of the entire population is susceptible to chronic infection with COVID-19. We also assume a stronger within-host selective pressure in chronically infected agents, compared to typical infected hosts.
To evaluate the model robustness and quantify the impact of the chronically infected hosts on the overall phylodynamics, we performed sensitivity analysis varying the percentage of chronically infected hosts between 0% to 5% (see Supplementary Section G).

### v3

@@ 4.2 Multi-scale phylodynamic simulation
This section describes processing layers (A), (C) and (D), highlighting their key dynamic relationships. Since the demographic characteristics of agents, including their social groups, are not updated during simulation, there is no dynamic modelling involved in layer (B), and this layer is described in S1 Text (Artificial agent-based population).
@@ 4.2.1 Phylogenetic layer (A)
The phylogenetic layer models the within-host evolution (mutation and selection), using pathogen genomes associated with each infected agent, and computes the fitness of circulating variants. An artificial genome contains 3,090 nucleotides (or 1,030 codons) which are associated with 20 known types of amino acids. The genome structure is partitioned into the spike region (100 codons) and the non-spike region (930 codons), with an overlapping region accounting for epitopes (75 codons, with 45 of these located in the spike region and 30 in the non-spike region). S1 Text (Genome structure) provides a detailed description of the genome structure, followed by S1 Text (Mutations) describing point mutations.
Pathogen fitness is defined in terms of its transmissibility, proportional to the corresponding basic or effective reproductive number ($R_{0}$ or $R_{\text{eff}}$). To quantify the individual fitness contributions of amino acids at each codon position, we employ a weight table, specifying weights $a_{i,j}$ of $N_{A}=20$ types of amino acids across $N=1,030$ codon positions, as detailed in S1 Text (Contribution to fitness). The overall fitness $K$ of a strain with genome $s$ is determined as the sum of these individual contributions:
@@
To re-iterate, the phylogenetic model distinguishes between two infected host categories: typical infected hosts and chronically infected hosts (S1 Text: Infected host categories). Consequently, there are different selective pressures driving the within-host evolution of pathogens in these categories (S1 Text: Intra-host (within-host) selective pressure). At each mutation step, we generate multiple candidate sequences, rank them by transmissibility, and focus on the group of top candidates. The size of this group indicates the strength of selective pressure. For example, selecting from a smaller group of top candidates implies a stronger selective pressure, while selecting from a larger group of top candidates implies a weaker selective pressure. When simulating selective pressure for chronically infected hosts, the size of the top group is assumed to be smaller, indicating a stronger selective pressure, compared to the selective pressure for typically infected hosts. We elaborate on this distinction and the limitation of these assumptions in Section 3.
We distinguish between mutation and substitution rates. While the mutation rate refers to the frequency of new mutations arising in a genome per unit of time, the substitution rate is the rate at which these new mutations are retained over time within the population [42, 43, 32]. In our case study, the mutation rate is a key input parameter, whereas the average substitution rate emerges as a simulation outcome. We use the attained substitution rate to calibrate the model, by matching the regression coefficient observed in empirical observations, illustrated in Fig 2B.
@@ 4.2.2 Epidemic layer (C)
As described in S1 Text (Multi-strain transmission model), disease is transmitted among agents that interact across different social contexts. The transmission process is simulated in discrete half-day time steps: “daytime” cycles during which agents interact in workplaces or educational settings (e.g., class, grade, school), and “nighttime” cycles during which agents interact in residential settings (e.g., household, household cluster, neighbourhood, and community). Each agent progresses through several health states: Susceptible, Infectious (asymptomatic or symptomatic), and Recovered, following the natural history of the disease.
At simulation cycle $n$, the infection probability $p_{i}(n)$ for a susceptible agent $i$ is determined across all its social contexts $g\in G_{i}$ (see S1 Text: Susceptible-infectious transition):
@@
where $q_{j\rightarrow i}(g)$ is the age-dependent interaction probability within $g$ (see S1 Text: Susceptible-infectious transition); $n_{j}$ denotes the infection onset time for agent $j$; the agent-specific function $f_{j}(n-n_{j})$ is the natural history of the disease, reflecting the infectivity of agent $j$ as its infection progresses (see S1 Text: Infectious-recovered and recovered-susceptible transitions); and $K(s_{j})$ represents the transmissibility of pathogen variant $s_{j}$, proportional to the corresponding basic reproductive number $R_{0}$ or effective reproductive number $R_{\text{eff}}$, i.e., $K(s_{j})$ is the fitness of genome $s_{j}$ carried by agent $j$, as defined by Eq 1.
Once susceptible agent $i$ becomes infected, it is possible to assign (i.e., identify) a specific infectious agent as the source of infection. This is simulated by weighted random sampling of an infection source from all potential infectious agents $j$ across all social contexts $g\in G_{i}(n)$ in which agent $i$ interacted during this cycle. Then the pathogen genome profile $s_{j}$ (carried by the identified infection source agent $j$) is copied to agent $i$ (see S1 Text: Susceptible-infectious transition).
Various interventions may change the infection probabilities across social contexts. Modelling non-pharmaceutical interventions (NPIs) is described in Section Non-pharmaceutical interventions.
@@ 4.2.3 Immunological layer (D)
The immunological layer simulates a vaccination rollout, including vaccination coverage, schedule and rates, given vaccine efficacy. Furthermore, the immuno-epidemiological model quantifies the agent immunity resulting from multiple immunity-boosting events (i.e., vaccinations and infections), as described in S1 Text (Vaccination). In this work, we broaden the concept of “hybrid immunity” to “compound immunity”, in order to capture non-linear immunity accumulation over various combinations of prior vaccinations and infections. Compound immunity may result from one or multiple vaccinations, one or multiple past infections, or a combination of both vaccination(s) and past infection(s) [28].
@@
where $m_{i}^{\textrm{c}}(n,r,s_{x})$ is the immunity against symptomatic infection induced by past immunological event $r$ (vaccination or infection); $H_{i}$ is the immunological history formed by past records $r$ up to cycle $n$; and $s_{x}$ is the genome carried by infectious agent $x$. In addition, the compound immunity wanes over time and depends on the genetic distance (see Eq 12 in S1 Text: Compound immunity against symptomatic infection). S1 Text (Compound immunity and waning effects) provides more details accounting for different vaccine efficacy components contributing to the compound immunity, and its effects on infection probabilities.
@@ 4.3.1 Hamming distance
Figs 2 and 3 trace the reference and pairwise Hamming distances for actual SARS-CoV-2 sequence data [15].
In order to trace Hamming distances for the simulated dynamics we followed slightly altered workflows, without applying the weekly windows for genome selection. When dealing with actual sequence data, these windows were needed to filter out sampling inconsistencies. Simulated data include pathogen genomes from all detected hosts, sampled on each simulation day, thus reducing sampling inconsistencies. The reference Hamming distance $\widehat{D}$ was computed against the simulated ancestral strain, constructed for each realisation (as described in Section 4). Figs 7, 911 trace the reference and pairwise Hamming distances for the simulated phylodynamics.
@@ 4.3.2 Statistical stationarity tests
ADF test. An ADF test detects non-stationarity in a time series [46]. We use the following null and alternative hypotheses to determine stationarity:
@@
We computed p-value from the ADF test and compared it against a chosen significance level (i.e., 0.05), with p-value smaller than the significance level rejecting $H_{0}$ and confirming stationarity. Results of ADF are shown in S1 Text (Stationarity of genomic diversity) and Fig 29.
CUSUM. A one-sided Cumulative Sum (CUSUM) control chart  [47, 48] can identify anomalies in the observed time series. In this study, we applied CUSUM on pairwise Hamming distance as shown in Fig 10.
@@
The high and low CUSUM values, $S^{X}$ and $S^{Y}$, are traced in Fig 10. We then applied a positive peak detection in $S^{X}$ and $S^{Y}$ (i.e., considering CUSUM value greater than 0), to detect the anomalies. Specifically, the peaks identified in $S^{X}$ indicate the emergence of a more transmissible variant, whereas the peaks identified in $S^{Y}$ indicate the dominance of a variant within the population. The number of dominant variants is equivalent to the total number of peaks detected in both $S^{X}$ and $S^{Y}$.
@@ Acknowledgement
Supplementary text also includes some additional results, sensitivity analysis and details on the computational complexity and implementation of PhASE TraCE.
@@ Appendix A Phylogenetic model
  • Genome structure represented by a nucleotide sequence with positions assigned to different functions.
  • Point mutations occurring at a rate that is averaged across the genome.
  • Pathogen fitness determined based on amino acids at specific regions of the genome.
  • Infected host categories, differentiating between typical infected hosts and chronically infected hosts.
  • Within-host (or intra-host) selective pressure generating advantageous pathogen mutations.
@@ A.1 Genome structure
Each nucleotide position in the simulated genome is represented by a number between 0 and 3. Every three nucleotides is interpreted as a codon (3,090 nucleotides correspond to 1,030 codons, see Fig 13), encoding a sequence of amino acids following the standard translation rules [49]. We annotate the 20 known amino acids by numerical labels from 0 to 19 instead of the conventional letters [49]. Although we model translation to an amino acid sequence, we do not look for coding regions or otherwise interpret or annotate the simulated sequence in the context of the real virus.
@@
The simulated genome is partitioned into two distinct regions (Fig 13): the “spike” region (consisting of 100 codons) and the “non-spike” region (consisting of 930 codons). This partitioning affects only two aspects of the simulation: (1) vaccine-related selective pressure is restricted to the spike region, and (2) mutations in the spike region have a greater chance of enhancing transmissibility (Section A.3). 75 codons (45 codons in the spike region and 30 codons in the non-spike region) are interpreted as epitopes relevant for modelling of immunity [50, 35]. 1,000 codons (including 45 that are also in epitopes) are interpreted as relevant to transmissibility. Section D.2.3 provides more details.
@@ A.2 Mutations
In each half-day simulation cycle, the total number of mutations ($L$) in each of the viral genomes within infected hosts is sampled from a Poisson distribution. The mean of the distribution is set based on an average mutation rate of 0.001 per nt per year across the genome, i.e., 3090×0.001/(365×2)0.004233090$\times$ 0.001/(365$\times$ 2$\approx$ 0.004233090 × 0.001 / ( 365 × 2 ) ≈ 0.00423. The average mutation rate was chosen within the reported range for SARS-CoV-2 of 0.000219 to 0.0012 per nt per year [36, 32]. The relationship between mutations and their contribution to fitness is described in Section A.3.
@@ A.3 Contribution to fitness
Individual mutations in the genome affect the overall viral fitness, resulting in growing transmissibility and higher reproductive number [17, 33, 34]. Obermeyer et al. [17] assessed the relative fitness of SARS-CoV-2 lineages (measured as the fold increase in $R$ over the ancestral strain) by linearly combining fitness contributions of individual amino acid substitutions, some of which were found to be fitness-increasing and recognised as spike mutations and non-spike mutations within the nucleocapsid and nonstructural proteins. Additionally, [34] identified a general distribution of amino acid selection at each position in SARS-CoV-2 sequences and highlighted positions that are prone to viral escape and mutations.
@@
Fig 14 shows a section of the weight table (the first 26 codon positions). The weight table generated for a simulation regulates the pathogen’s evolution. In the absence of other pressures, one would expect substitutions that increase fitness to be preferentially fixed as a simulation progresses, leading to a gradual increase in transmissibility. Using the weight table, we calculate the overall fitness $K$ of a genome $s$ by combining the individual contributions of the present amino acids:
@@
where $a_{i,s(i)}$ is the fitness contribution of amino acid $j=s(i)$ located at codon position $i$, with $j\in[0,N_{A}-1]$. We note that $a_{i,s(i)}$ can be zero, positive, or negative, representing mutations of neutral, advantageous, or disadvantageous contribution to pathogen fitness. We note that Eq 7 and Eq 1 in the main manuscript are equivalent.
@@
Prior to simulating the COVID-19 case study, we generated a genome for the ancestral SARS-CoV-2 strain sampled such that its viral fitness, calculated using Eq 7 and the weight table, fell between 2.65 and 2.85, matching the transmissibility range of the ancestral SARS-CoV-2 strain modelled in our previous work [6]. We used this ancestral genome as a starting point from which other variants originate, thus assuming a shared ancestry for all variants [52, 53]. To optimise memory usage, given the high number of simulated infections, each strain is represented by a list of mutations against the ancestral genome instead of a full-length genome.
@@ A.4 Infected host categories
Host factors, including individual variation in immune response, are known to strongly affect epidemic progression [54, 55]. In particular, chronic carriage of SARS-CoV-2 by immunocompromised hosts is thought to have had a key role in the emergence of new viral variants during the pandemic [32]. In a host with compromised or suppressed immune function, the virus is subjected to different selective pressures and within-host dynamics. A host immune system that is not able to effectively clear the virus is likely to result in a longer infection duration, a higher viral load within the host, and a greater diversity of viral sub-populations [32]. These conditions may facilitate more substitutions and larger increases in viral fitness (saltational evolution) compared to the evolution of the virus over the same interval of time when transmitted by a chain of typical hosts and subjected to transmission bottlenecks and competent immune systems. In the case of SARS-CoV-2, chronically infected hosts can have a prolonged recovery period of several months, compared to one to two weeks for typical infected hosts [29, 30].
To examine the possible role of chronic infections on the pathogen evolution, we differentiate between two infected host categories: chronically infected hosts and the typical infected hosts. The main distinctions include the following aspects: (i) chronically infected hosts are assigned a different natural history model with a significantly longer recovery period, randomly sampled from a uniform distribution ranging from 60 to 370 days reported in many studies (Fig 15[30, 31, 56], and (ii) chronically infected hosts may have a higher intra-host selective pressure (detailed in Section G and Fig 32).
Prior study [30] reported that chronically infected hosts make up a small fraction of the infected hosts (between 0.1 and 0.5% of COVID-19 infections), and only 32% of these hosts may experience strong positive intra-host selection. Considering the high attack rate of the COVID-19, we assume that 0.1% of the entire population is susceptible to chronic infection with COVID-19. We also assume a stronger within-host selective pressure in chronically infected agents, compared to typical infected hosts.
To evaluate the model robustness and quantify the impact of the chronically infected hosts on the overall phylodynamics, we performed sensitivity analysis varying the percentage of chronically infected hosts between 0% to 5% (see Section G, and Fig 31).